OpenStudy (anonymous):

How to SOLVE: medals can be given if it satisfies u Integrate the following indefinite integral: (sin7x)^12 * (cos7x)^3 Hint: sin^2 + cos^2 = 1

OpenStudy (anonymous):

OK, you're NOT gonna want to do anything with it but: { 1/22364160 } * { 13*\sin(105x) - 135*\sin(91x) + 585*\sin(77x) - 1235*\sin(63x) + 585*\sin(49x) + 3861*\sin(35x) - 12155*\sin(21x) + 19305*\sin(7x) } + C

OpenStudy (anonymous):

check it yourself

OpenStudy (anonymous):

any objections

OpenStudy (anonymous):

This integral is why you memorize the reduction formulas. http://en.wikipedia.org/wiki/Integration_by_reduction_formulae

OpenStudy (anonymous):

it was just how to solve

OpenStudy (anonymous):

initially hows that

OpenStudy (anonymous):

You could also do it by painfully redefining each\[\sin^2x=1-\cos^2x\]and its reverse alternative. If you run into a problem, I think dividing all sides by \(\cos^2x\) for an easy \(\sec^2x\to\tan x\) relation would be the best bet.

OpenStudy (lgbasallote):

how about turning \[(\cos 7x)^3 \implies (1 - \sin^2 7x)(\cos 7x)\]

OpenStudy (lgbasallote):

\[\int (\sin^{12} 7x - \sin^{14} 7x)(\cos x)dx\]

OpenStudy (lgbasallote):

now it's u-sub

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then

OpenStudy (lgbasallote):

lol should i have to complete it =))

OpenStudy (anonymous):

no medal me if my answer satisfies u

OpenStudy (anonymous):

\[ (sin7x)^{12} * (cos7x)^3=(sin7x)^{12} * (cos7x)^2 (cos7x) = (sin7x)^{12} (1- (sin7x))^2 (cos7x) \]

OpenStudy (anonymous):

u= sin( 7x) and the rest is easy.

OpenStudy (lgbasallote):

\[\frac 17\int \sin^{12} u)(\cos u) du - \frac 17 \int (\sin^{14} u )(\cos u du\]

OpenStudy (anonymous):

You probably know this, but use \(\text{\\}\) in \(\TeX\) if you want a line break.

OpenStudy (anonymous):

as u r an retired professer

OpenStudy (lgbasallote):

\[\frac{\sin^{13}u}{91} - \frac{\sin^{15} u}{105}\]

OpenStudy (lgbasallote):

+C

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

hmmmmmmmmmm

OpenStudy (lgbasallote):

happy now lol

OpenStudy (anonymous):

@eliassaab I saw u r website it helped me a lot

OpenStudy (anonymous):

Thanks @Rohangrr

OpenStudy (anonymous):

@eliassaab you wrote cos7x=1-sin7x...is it true?

OpenStudy (anonymous):

yes its

OpenStudy (lgbasallote):

so roh...

OpenStudy (anonymous):

Sorry \[ 1 - \sin^2(7x) \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

lgb

OpenStudy (lgbasallote):

lol still not happy?

OpenStudy (anonymous):

@Rohangrr give me a medal man!!

OpenStudy (anonymous):

why soooooo

OpenStudy (lgbasallote):

it's obviously right

OpenStudy (anonymous):

elisaab deserved the medal from my side

OpenStudy (anonymous):

i found a big mistake if not ur ans will be wrong

OpenStudy (anonymous):

okay

OpenStudy (lgbasallote):

lol whatever...what i want is you to admit i was right

OpenStudy (anonymous):

wait got it

OpenStudy (anonymous):

\[ (sin7x)^{12} * (cos7x)^3=(sin7x)^{12} * (cos7x)^2 (cos7x) =\\ (sin7x)^{12} (1- (sin7x)^2) (cos7x) \]

OpenStudy (anonymous):

well I am running behind the answer

OpenStudy (anonymous):

or running behind the answer

OpenStudy (anonymous):

@lgbasallote i think u made a mistake \[(\sin 7x)^{12} \] not \[\sin^{12} (7x)\]

OpenStudy (lgbasallote):

\[(\sin 7x)^12 \implies \sin^{12} 7x \cdots\]

OpenStudy (lgbasallote):

^you now what i mean lol

OpenStudy (anonymous):

I dislike the notation as well, since it isn't internally consistent with\[\arcsin x=\sin^{-1}x\neq\frac1{\sin x}\]

OpenStudy (anonymous):

@lgbasallote sorry Oh, what did I say my mistake ;)

OpenStudy (lgbasallote):

i made no mistake :) @Rohangrr just refuses to admit im right haha